OpenStudy (anonymous):

Which of the following is the equation of a line perpendicular to the line y=-3/2 x+4, passing through the point (3,9)? -2x-3y=-21 2x-3y=21 2x+3y=21 -2x+3y=21

OpenStudy (freckles):

can you identify the slope of the line y=-3/2 x+4 ?

OpenStudy (freckles):

well the slope of y=mx+b is m so your slope is...?

OpenStudy (freckles):

Once you identify the slope of y=-3/2 x+4 we can find the slope of the perpendicular line For my line y=mx+b, a perpendicular line to that could be given by y=-1/m x+c . You find what c needs to be by entering in your point in which this line should go through .

OpenStudy (freckles):

I'm going to take off now. Sleep time.

OpenStudy (anonymous):

Find the slope of this line. The slope of your perpendicular will be the negative reciprocal of this.